How to find equation of lines from pair of straight lines?

Grade:12

1 Answers

Vikas TU
6 years ago
 Attempting to illuminate x2−3xy+ 2y2 
=x2−2xy−xy+2y2 
=x(x−2y)−y(x−2y) 
=(x−y)(x−2y) 
Presently other direct and consistent terms will be expected to an and b. 
(x−y+a)(x−2y+b)=x2−3xy+2y2+x−2 
or,bx+ax+ab−by−2ay=x−2 
or,(a+b)x+ab−(2a+b)y=x−2 
a+b=1(i) 
2a+b=0 (ii) 
Subtracting (i) from (ii) 
a=−1 
At that point, b=2 
In this way, 
(x−y−1)(x−2y+2)=0 
Either, 
x−y−1=0 
or, on the other hand, 
x−2y+2=0 
On the off chance that both are zero all the while we can discover arrangement, 
y−3=0 
y=3 andx=4